what will be the width of a rectangle with the length of 3/4 and the area of 1/8

Answer by hamsanash1981@gmail.com(151)

length = 3/4
area = 1/8
breath =x
l*b = area
3/4*x = 1/8
x = 1/8*4/3
x = 1/6
therefore breath = 1/6